OpenStudy (anonymous):

solve by the subst. method : y=3x-12 2x-3y=-13

OpenStudy (day_2010):

Wrong section :) You will get better help in the math section.

OpenStudy (anonymous):

y=3x-12......1 2x-3y=-13......2 substitute equation y into equation 2 2x-3[3x12]=-13 2x-9x+36=-13 collect like terms -7x=-13-36 -7x=-49 divide through by-7 -7x/-7=-49/-7 x=7 substitute x into equation 1 y=3[7]-12 y=21-12 y=9 therefore,x=7 and y=9

OpenStudy (day_2010):

Hi, Can you close this question @breeskee03. Obviously, you acquired your 'answer' on the writing section. Thanks :)